OpenStudy (anonymous):

If (x + 1)(x - 3) = 12, then which of the following statements is true? x + 1 = 0 or x - 3 = 0 x + 1 = 12 or x - 3 = 12 x - 5 = 0 or x + 3 = 0

OpenStudy (anonymous):

x-5 and x+3 is true.

OpenStudy (anonymous):

(x + 1)(x - 3) = 12 x^2 -3x +x -3 = 12 x^2 -2x -3 = 12 x^2 -2x -15 = 0; x = -3 & x = 5